### v1

# Logarithmic Sobolev inequality in manifolds with nonnegative curvature via the ABP method

Abstract.
In this paper, we employ the ABP method developed by Brendle [5] to establish the optimal $L^{p}$-logarithmic-Sobolev inequality on manifolds with nonnegative Ricci curvature, as well as a sharp $L^{2}$-logarithmic-Sobolev inequality for submanifolds in manifolds with nonnegative sectional curvature. The sharp constants in both inequalities depend on the asymptotic volume ratio of the ambient manifold.
1. Introduction
The classical sharp logarithmic Sobolev inequality on $\mathbb{R}^{n}$, proved by Gross [14], establishes that
$\int_{\mathbb{R}^{n}}f^{2}\log f^{2}\,d\gamma-\Big(\int_{\mathbb{R}^{n}}f^{2}\,d\gamma\Big)\log\Big(\int_{\mathbb{R}^{n}}f^{2}\,d\gamma\Big)\leq 2\int_{\mathbb{R}^{n}}|Df|^{2}\,d\gamma,$
where $d\gamma=(2\pi)^{-\frac{n}{2}}e^{-\frac{|x|^{2}}{2}}\,dx$ is the standard Gaussian probability measure and $f\in C_{c}^{\infty}(\mathbb{R}^{n})$. This is equivalent to the optimal $L^{2}$-logarithmic Sobolev inequality established by Weissler [20]:
$\int_{\mathbb{R}^{n}}f^{2}\log f^{2}\,dx\leq\frac{n}{2}\log\Big(\frac{2}{\pi en}\int_{\mathbb{R}^{n}}|Df|^{2}\,dx\Big),$
valid for $f\in C_{c}^{\infty}(\mathbb{R}^{n})$ satisfying $\int_{\mathbb{R}^{n}}f^{2}\,dx=1$.
The following Euclidean $L^{p}$-logarithmic Sobolev inequality, established by Beckner [2] for $p=1$, Del Pino and Dolbeault [11] for $1<p<n$ and Gentil [13] for general $p>1$, asserts that for $f\in C_{c}^{\infty}(\mathbb{R}^{n})$ with $\int_{\mathbb{R}^{n}}f^{p}\,dx=1$,
$\int_{\mathbb{R}^{n}}|f|^{p}\log|f|^{p}\,dx\leq\frac{n}{p}\log\Big(\mathrm{L}(n,p)\int_{\mathbb{R}^{n}}|Df|^{p}\,dx\Big),$
where the sharp constant $\mathrm{L}(n,p)$ is given by
$\mathrm{L}(n,p)=\left\{\begin{array}[]{ll}n^{-1}\omega_{n}^{-\frac{1}{n}}&\text{if }p=1,\\ \frac{p}{n}\Big(\frac{p-1}{e}\Big)^{p-1}\Big(\omega_{n}\Gamma\big(\frac{n}{q}+1\big)\Big)^{-\frac{p}{n}}&\text{if }p>1.\end{array}\right$.
with $\Gamma$ the Gamma function, $q=\frac{p}{p-1}$ the conjugate of $p$, and $\omega_{n}=\frac{\pi^{\frac{n}{2}}}{\Gamma(\frac{n}{2}+1)}$ the volume of the unit ball in $\mathbb{R}^{n}$.
Recently, by generalizing the optimal transport approach developed by Cordero-Erausquin, Nazaret and Villani [10], Kristály [16] obtained the optimal $L^{p}$-logarithmic Sobolev inequality for manifolds with nonnegative Ricci curvature. In this paper, inspired by the ABP method developed by Brendle [5], our first goal is to provide a ABP-based proof of this inequality.
Let $(M^{n},g)$, with $n\geq 2$, be a complete noncompact manifold with nonnegative Ricci curvature. Denote the asymptotic volume ratio of $M$ by
$\theta:=\lim_{r\to\infty}\frac{|B_{o}(r)|}{\omega_{n}r^{n}},$
where $B_{o}(r)$ is the metric ball centered at $o\in M$ with radius $r$. The Bishop-Gromov volume comparison theorem implies that $\theta\leq 1$. If $\theta>0$, then we say that $M$ has Euclidean volume growth.
The sharp $L^{1}$-logarithmic Sobolev inequality for manifolds with nonnegative Ricci curvature, proved using the ABP method developed by Brendle [5], reads as follows:
Theorem 1.1.
Let $(M^{n},g)$, with $n\geq 2$, be a complete noncompact Riemannian manifold with nonnegative Ricci curvature and Euclidean volume growth. Let $\Omega$ be a compact domain in $M$ with boundary $\partial\Omega$, and let $f$ be a positive smooth function in $\Omega\setminus\partial\Omega$. Then
$\int_{\Omega}f\log f\,d\mathrm{vol}\leq n\log\bigg(n^{-1}\omega_{n}^{-\frac{1}{n}}\theta^{-\frac{1}{n}}\Big(\int_{\Omega}|Df|\,d\mathrm{vol}+\int_{\partial\Omega}f\,da\Big)\bigg),$
where $\theta$ denotes the asymptotic volume ratio of $M$.
The sharp $L^{p}(p>1)$ logarithmic Sobolev inequality for manifolds with nonnegative Ricci curvature, which we established based on the ABP method developed by Brendle [5], reads as follows:
Theorem 1.2.
Let $(M^{n},g)$, with $n\geq 2$, be a complete noncompact Riemannian manifold with nonnegative Ricci curvature and Euclidean volume growth. Let $\Omega$ be a compact domain in $M$ with boundary $\partial\Omega$, and let $f$ be a positive smooth function in $\Omega\setminus\partial\Omega$ that vanishes on $\partial\Omega$. For $p>1$, we have
$\int_{\Omega}f^{p}\log f^{p}\,d\mathrm{vol}\leq\frac{n}{p}\log\bigg(\mathrm{L}(n,p)\theta^{-\frac{p}{n}}\int_{\Omega}|Df|^{p}\,d\mathrm{vol}\bigg),$
where $\theta$ denotes the asymptotic volume ratio of $M$.
We now turn to the logarithmic Sobolev inequality for submanifolds. In [12], Ecker derived such an inequality valid for submanifolds in Euclidean space. Later in 2020, Brendle [4] established a sharp logarithmic Sobolev inequality for compact submanifolds in Euclidean space without boundary, employing the ABP techniques developed in [3]. In 2021, Yi and Zheng [22] proved a logarithmic Sobolev inequality for compact boundaryless submanifolds in manifolds with nonnegative sectional curvature. Recently, by developing a noncompact modification of Brendle’s ABP method, Wang, Xia and Zhang [18] obtained an optimal logarithmic Sobolev inequality for non-compact self-shrinkers in Euclidean space. Independently, Balogh and Kristály [1] established the sharp logarithmic Sobolev inequality for submanifolds in Euclidean space using an optimal transport approach. Inspired by [4, 5, 22], our second goal herein is to provide an ABP-based proof of the optimal logarithmic Sobolev inequality for submanifolds in manifolds with nonnegative sectional curvature.
The following theorem extends Theorem 1.1 of [1] to submanifolds in manifolds with nonnegative sectional curvature.
Theorem 1.3.
Let $(M^{n+m},g)$, with $n\geq 2$ and $m\geq 1$, be a complete noncompact Riemannian manifold with nonnegative sectional curvature and Euclidean volume growth. Let $\Sigma^{n}$ be a compact submanifold of $M$, possibly with boundary, and let $f$ be a positive smooth function in $\Sigma\setminus\partial\Sigma$ that vanishes on $\partial\Sigma$. Then
$\int_{\Sigma}f^{2}\log f^{2}\leq\frac{n}{2}\log\Big(\frac{2}{\pi en}\theta^{-\frac{n}{2}}\int_{\Sigma}|D^{\Sigma}f|^{2}+\frac{1}{4}f^{2}|H|^{2}\,d\mathrm{vol}\Big).$
where $H$ is the mean curvature vector of $\Sigma$ and $\theta$ denotes the asymptotic volume ratio of $M$.
The ABP method, originating from the work of Alexandrov, Bakelman, and Pucci, is fundamental in PDE theory. Cabré was the first to use the ABP method to provide a simple proof of the classical isoperimetric inequality in $\mathbb{R}^{n}$ around 1996 (see [7]). Recently, by generalizing the ABP technique developed by Cabré [8], Brendle [3] established a sharp isoperimetric inequality for minimal submanifolds in Euclidean space of codimension at most $2$. For further details regarding the ABP method, we refer the reader to [6, 9, 15, 17, 19, 21] and the references therein.
This paper is organized as follows: Section 2 is devoted to the ABP-based proof of Theorem 1.1 and Theorem 1.2. In Section 3, we give an ABP-based proof of Theorem 1.3.
2. The case of domains
Throughout this section, we assume that $(M^{n},g)$, for $n\geq 2$, is a complete noncompact manifold with nonnegative Ricci curvature, and we fix a base point $o\in M$. Let $\Omega\subset M$ be a compact domain with smooth boundary $\partial\Omega$. Let $p\geq 1$ and let $f$ be a smooth positive function in $\Omega\setminus\partial\Omega$. In the case $p>1$, we assume that $f$ vanishes on $\partial\Omega$.
without loss of generality, we assume hereafter that $\Omega$ is connected. Let $\alpha>0$ be a given parameter and $r_{o}=d(o,\Omega)$. By sacling, we may assume that
(1) $\left\{\begin{aligned} &\int_{\Omega}f^{p}\log f^{p}\,d\text{vol}-\frac{1}{\alpha}\int_{\Omega}|Df|^{p}\,d\text{vol}=\frac{1}{\alpha}\int_{\partial\Omega}f\,da,&p=1,\\ &\int_{\Omega}f^{p}\log f^{p}d\text{vol}-\frac{1}{\alpha}\int_{\Omega}|Df|^{p}d\text{vol}=0,&p>1.\end{aligned}\right$.
Since $\Omega$ is connected, we can find a solution of the following Neumann boundary problem
(2) $\left\{\begin{aligned} &\mathrm{div}(f^{p}Du)=f^{p}\log f^{p}-\frac{1}{\alpha}|Df|^{p},&\text{ in }\Omega\setminus\partial\Omega,\\ &\langle Du,\nu\rangle=\frac{r^{p-1}}{\alpha},&\text{ on }\partial\Omega,\end{aligned}\right$.
where $\nu$ is the outward unit normal vector field and $r>1+\alpha r_{o}$ is an arbitrary constant. By standard elliptic regularity theory, we know that $u\in C^{2,\gamma}$ for each $0<\gamma<1$.
As in [5], we set $U:=\{x\in\Omega\setminus\partial\Omega:|Du(x)|<\frac{r^{p-1}}{\alpha}\}$ and
$A_{r}=\{\bar{x}\in U:ru(x)+\frac{1}{2}{d}(x,\exp_{\bar{x}}(rDu(\bar{x})))^{2}\geq ru(\bar{x})+\frac{1}{2}r^{2}|Du(\bar{x})|^{2},\ \forall x\in\Omega\}.$
Moreover, we define the transport map $\Phi_{r}:\Omega\to M$ by
$\Phi_{r}(x)=\exp_{x}(rDu(x)),$
for all $x\in\Omega$. Note that the map $\Phi_{r}$ is of class $C^{1,\gamma}$ for each $0<\gamma<1$.
The proofs of the following two lemmas are owed to Brendle, so we omit them here.
Lemma 2.1.
The set
$\Omega_{r}:=\{\bar{x}\in M:d(x,\bar{x})<\frac{r^{p}}{\alpha},\quad\forall x\in\Omega\}$
is contained in $\Phi_{r}(A_{r})$.
Proof.
See Lemma 2.2 in [5]. ∎
Lemma 2.2.
The Jacobian determinant of $\Phi_{r}$ satisfies
$|\det D\Phi_{r}(x)|\leq(1+r\frac{\Delta u(x)}{n})^{n}$
for all $x\in A_{r}$.
Proof.
See Corollary 2.6 in [5]. ∎
Lemma 2.3.
The Jacobian determinant of $\Phi_{r}$ satisfies
$\left\{\begin{aligned} &|\det D\Phi_{r}(x)|\leq r^{n}e^{\frac{n}{r}-n}f(x),&p=1,\\ &e^{-\frac{p}{q}\alpha^{\frac{q}{p}}\frac{|d(x,\Phi_{r}(x))|^{q}}{r^{q}}}|\det D\Phi_{r}(x)|\leq r^{n}e^{\frac{n}{r}-n}f(x)^{p},&p>1.\end{aligned}\right$.
for all $x\in A_{r}$.
Proof.
By (2), we have
$\mathrm{div}(f^{p}Du)$ $=f^{p}\Delta u+p\langle f^{p-1}Df,Du\rangle$
$=f^{p}\log f^{p}-\frac{1}{\alpha}|Df|^{p}.$
In the case $p=1$, using $|Du(x)|<\frac{1}{\alpha}$ and the Cauchy-Schwarz inequality, one has
$\Delta u$ $=\log f-\frac{1}{\alpha}\frac{|Df|}{f}-\langle\frac{Df}{f},Du\rangle$
$\leq\log f.$
By Lemma 2.2 and the elementary inequality $\lambda\leq e^{\lambda-1}$, it follows that
$|\det D\Phi_{r}(x)|$ $\leq\Big(\frac{1}{r}+\frac{\Delta u(x)}{n}\Big)^{n}r^{n}$
$\leq r^{n}e^{\frac{n}{r}+\Delta u-n}$
$=r^{n}e^{\frac{n}{r}-n}f(x).$
For $p>1$, through the Young’s inequality, one has
$\Delta u$ $=\log f^{p}-\frac{|Df|^{p}}{\alpha f^{p}}-p\langle\frac{Df}{f},Du\rangle$
$\leq\log f^{p}+\frac{p}{q}\alpha^{\frac{q}{p}}|Du|^{q}.$
Similarly, one can get
$|\det D\Phi_{r}(x)|$ $\leq\Big(\frac{1}{r}+\frac{\Delta u(x)}{n}\Big)^{n}r^{n}$
$\leq r^{n}e^{\frac{n}{r}+\Delta u(x)-n}$
$\leq r^{n}e^{\frac{n}{r}-n}e^{\frac{p}{q}\alpha^{\frac{q}{p}}|Du(x)|^{q}}f(x)^{p},$
which implies
$e^{-\frac{p}{q}\alpha^{\frac{q}{p}}\frac{|d(x,\Phi_{r}(x))|^{q}}{r^{q}}}|\det D\Phi_{r}(x)|\leq r^{n}e^{\frac{n}{r}-n}f(x)^{p}.$
Proof of Theorem 1.1.
Using the triangle inequality, Lemma 2.1 and Lemma 2.3, the formula for change of variables in multiple integrals implies
$\theta\omega_{n}\big(\frac{r}{\alpha}-r_{o}\big)^{n}$ $\leq\Big|B_{o}\big(\frac{r}{\alpha}-r_{o}\big)\Big|$
$\leq|\{\bar{x}\in M:d(x,\bar{x})<\frac{r}{\alpha},\,\forall x\in\Omega\}|$
$\leq\int_{A_{r}}|D\Phi_{r}(x)|\,d\text{vol}(x)$
$\leq r^{n}e^{\frac{n}{r}-n}\int_{\Omega}f(x)\,d\text{vol}(x)$
for all $r>1+\alpha r_{o}$. Dividing the above formula by $r^{n}$ and let $r\to\infty$, we obtain
$\theta\omega_{n}\alpha^{-n}\leq e^{-n}\int_{\Omega}f\,d\text{vol},$
which implies
$n+\log(\theta\omega_{n}\alpha^{-n})\leq\log\bigg(\int_{\Omega}f\,d\text{vol}\bigg).$
Combining this inequality with (1) yields the following scale-invariant inequality:
(3) $\int_{\Omega}f(\log f+n+\log(\theta\omega_{n}\alpha^{-n}))\,d\text{vol}-\frac{1}{\alpha}\int_{\Omega}|Df|\,d\text{vol}-\frac{1}{\alpha}\int_{\partial\Omega}f\,da$
$\leq\Big(\int_{\Omega}f^{p}\,d\mathrm{vol}\Big)\log\Big(\int_{\Omega}f^{p}\,d\mathrm{vol}\Big)$
Let $\int_{\Omega}f\,d\text{vol}=1$, we have
$\int_{\Omega}f\log f\,d\text{vol}\leq-n-\log(\theta\omega_{n}\alpha^{-n})+\frac{1}{\alpha}\Big(\int_{\Omega}|Df|\,d\text{vol}+\int_{\partial\Omega}f\,da\Big)$
By minimizing the right hand side with respect to $\alpha>0$, one can get
$\alpha=\frac{1}{n}\Big(\int_{\Omega}|Df|\,d\text{vol}+\int_{\partial\Omega}f\,da\Big).$
This implies
$\int_{\Omega}f\log f\,d\text{vol}\leq n\log\bigg(n^{-1}\omega_{n}^{-\frac{1}{n}}\theta^{-\frac{1}{n}}\Big(\int_{\Omega}|Df|\,d\text{vol}+\int_{\partial\Omega}f\,da\Big)\bigg).$
Proof of Theorem 1.2.
Let $p>1$. By the triangle inequality, Lemma 2.1 and Lemma 2.3, we have
$\int_{\Omega_{r}}e^{-\frac{p}{q}\alpha^{\frac{q}{p}}\frac{(r_{o}+d(o,\bar{x}))^{q}}{r^{q}}}\,d\mathrm{vol}(\bar{x})$
$\leq\int_{\Omega}e^{-\frac{p}{q}\alpha^{\frac{q}{p}}\frac{(d(x,\Phi_{r}(x)))^{q}}{r^{q}}}|\det D\Phi_{r}(x)|1_{A_{r}}(x)\,d\mathrm{vol}(x)$
$\leq r^{n}e^{\frac{n}{r}-n}\int_{\Omega}f(x)^{p}\,d\mathrm{vol}(x).$
Note that $B_{o}\Big(\frac{r^{p}}{\alpha}-r_{o}\Big)\subset\Omega_{r}$, the co-area formula implies
$\int_{0}^{\frac{r^{p}}{\alpha}-r_{o}}e^{-\frac{p}{q}\alpha^{\frac{q}{p}}\frac{(r_{o}+\rho)^{q}}{r^{q}}}n\omega_{n}\theta\rho^{n-1}\,d\rho\leq r^{n}e^{\frac{n}{r}-n}\int_{\Omega}f(x)^{p}\,d\mathrm{vol}(x).$
Dividing the above formula by $r^{n}$ and let $r\to\infty$, it follows from (A.1) in [1] that
$\theta e^{n}\bigg(\Big(\frac{p}{q}\alpha^{\frac{q}{p}}\Big)^{-\frac{n}{q}}\omega_{n}\Gamma\Big(\frac{n}{q}+1\Big)\bigg)$
$=\lim_{r\to\infty}r^{-n}e^{n-\frac{n}{r}}\int_{0}^{\frac{r^{p}}{\alpha}-r_{o}}e^{-\frac{p}{q}\alpha^{\frac{q}{p}}\frac{(r_{o}+\rho)^{q}}{r^{q}}}n\omega_{n}\theta\rho^{n-1}\,d\rho$
$\leq\int_{\Omega}f(x)^{p}\,d\mathrm{vol}(x).$
Combining this inequality with (1) gives
$\int_{\Omega}f^{p}\bigg(\log f^{p}+n+\log\theta-\frac{n}{q}\log\Big(\frac{p}{q}\alpha^{\frac{q}{p}}\Big)+\log\Big(\omega_{n}\Gamma\Big(\frac{n}{q}+1\Big)\Big)\bigg)\,d\mathrm{vol}$
$\leq\frac{1}{\alpha}\int_{\Omega}|Df|^{p}\,d\mathrm{vol}+\Big(\int_{\Omega}f^{p}\,d\mathrm{vol}\Big)\log\Big(\int_{\Omega}f^{p}\,d\mathrm{vol}\Big).$
Let $\int_{\Omega}f^{p}\,d\text{vol}=1$, we have
$\int_{\Omega}f^{p}\log f^{p}\,d\mathrm{vol}$
$\leq-n-\log\theta+\frac{n}{q}\log\Big(\frac{p}{q}\alpha^{\frac{q}{p}}\Big)-\log\Big(\omega_{n}\Gamma\Big(\frac{n}{q}+1\Big)\Big)\bigg)+\frac{1}{\alpha}\int_{\Omega}|Df|^{p}\,d\mathrm{vol}.$
Minimizing the right-hand side with respect to $\alpha>0$ yields
$\alpha=\frac{p}{n}\int_{\Omega}|Df|^{p}\,d\mathrm{vol},$
which implies
$\int_{\Omega}f^{p}\log f^{p}\,d\mathrm{vol}\leq\frac{n}{p}\log\bigg(\mathrm{L}(n,p)\theta^{-\frac{p}{n}}\int_{\Omega}|Df|^{p}\,d\mathrm{vol}\bigg).$
3. The case of submanifolds
In this section, we assume that $(M^{n+m},g)$, with $n\geq 2$ and $m\geq 1$, is a complete noncompact manifold with nonnegative sectional curvature, and we fix a base point $o\in M$. Let $\Sigma\subset M$ be a compact submanifold of dimension $n$, possibly with boundary. Let $p=2$ and let $f$ be a smooth positive function defined in $\Sigma\setminus\partial\Sigma$ that vanishes on $\partial\Sigma$. Let $\bar{D}$ denote the the Levi-Civita connection of $M$, and let $D^{\Sigma}$ denote the induced connection on $\Sigma$. The second fundamental form $B$ of $\Sigma$ is defined by
$\langle B(X,Y),V\rangle=\langle\bar{D}_{X}Y,V\rangle,$
where $X$ and $Y$ are tangent vector fields on $\Sigma$, and $V$ is a normal vector field along $\Sigma$. In particular, the mean curvature vector $H$ is defined as the trace of the second fundamental form $B$.
It suffies to prove Theorem 1.3 for the case where $\Sigma$ is connected. Let $\alpha>0$ be a given parameter and $r_{o}=d(o,\Sigma)$. By scaling, we may assume that
(4) $\int_{\Sigma}f^{2}\log f^{2}\,d\mathrm{vol}-\frac{1}{\alpha}\int_{\Sigma}|D^{\Sigma}f|^{2}\,d\mathrm{vol}-\frac{1}{4\alpha}\int_{\Sigma}f^{2}|H|^{2}\,d\mathrm{vol}=0.$
By the connectedness of $\Sigma$, there exists a solution of the following Neumann boundary problem
(5) $\left\{\begin{aligned} &\mathrm{div}(f^{2}D^{\Sigma}u)=f^{2}\log f^{2}-\frac{1}{\alpha}|D^{\Sigma}f|^{2}-\frac{1}{4\alpha}f^{2}|H|^{2},&\text{ in }\Sigma\setminus\partial\Sigma,\\ &\langle D^{\Sigma}u,\nu\rangle=\frac{r}{\alpha},&\text{ on }\partial\Sigma,\end{aligned}\right$.
where $\nu$ is the outward unit normal vector field of $\partial\Sigma$ with respect to $\Sigma$ and $r>1+\alpha r_{o}$ is an arbitrarily constant. By standard elliptic regularity theory, we know that $u\in C^{2,\gamma}$ for each $0<\gamma<1$.
As in [5], we set $U:=\{x\in\Sigma\setminus\partial\Sigma:|Du(x)|<\frac{r}{\alpha}\}$ and
$E:=\{(x,y):x\in U,y\in T^{\perp}_{x}\Sigma,|D^{\Sigma}u(x)|^{2}+|y|^{2}<\frac{r^{2}}{\alpha^{2}}\}.$
Let $A_{r}$ be the set of all points $(\bar{x},\bar{y})\in E$ satisfying
$ru(x)+\frac{1}{2}{d}(x,\exp_{\bar{x}}(rD^{\Sigma}u(\bar{x})+r\bar{y}))^{2}\geq ru(\bar{x})+\frac{1}{2}r^{2}(|D^{\Sigma}u(\bar{x})|^{2}+|\bar{y}|^{2})$
for all $x\in\Sigma$. Moreover, we define the transport map $\Phi_{r}:T^{\perp}\Sigma\to M$ by
$\Phi_{r}(x,y)=\exp_{x}(rD^{\Sigma}u(x)+ry)$
for all $x\in\Sigma$ and $y\in T^{\perp}_{x}\Sigma$. Note that the map $\Phi_{r}$ is of class $C^{1,\gamma}$ for each $0<\gamma<1$.
The proofs of the following two lemmas are owed to Brendle, so we omit them here.
Lemma 3.1.
The set
$\Sigma_{r}:=\{\xi\in M:d(x,\xi)<\frac{r^{2}}{\alpha}\text{ for all }x\in\Sigma\}$
is contained in $\Phi_{r}(A_{r})$.
Proof.
See Lemma 4.2 in [5]. ∎
Lemma 3.2.
The Jacobian determinant of $\Phi_{r}$ satisfies
$|\det D\Phi_{r}(x,y)|\leq r^{n+m}\Big(\frac{1}{r}+\frac{\Delta_{\Sigma}u(x)-\langle H(x),y\rangle}{n}\Big)^{n}$
for all $(x,y)\in A_{r}$.
Proof.
See Corollary 4.7 in [5]. ∎
Lemma 3.3.
The Jacobian determinant of $\Phi_{r}$ satisfies
$e^{-\alpha\frac{|d(x,\Phi_{r}(x,y))|^{2}}{r^{2}}}|\det D\Phi_{r}(x,y)|\leq r^{n+m}e^{\frac{n}{r}-n}f(x)^{2}$
for all $(x,y)\in A_{r}$.
Proof.
By (5), we have
$\mathrm{div}(f^{2}D^{\Sigma}u)$ $=f^{2}\Delta_{\Sigma}u+2\langle fD^{\Sigma}f,D^{\Sigma}u\rangle$
$=f^{2}\log f^{2}-\frac{1}{\alpha}|D^{\Sigma}f|^{2}-\frac{1}{4\alpha}f^{2}|H|^{2}.$
Using the Cauchy-Schwarz inequality, one can derive that
$\Delta_{\Sigma}u(x)-\langle H(x),y\rangle$
$=\log f(x)^{2}-\frac{|D^{\Sigma}f(x)|^{2}}{\alpha f(x)^{2}}-\frac{1}{4\alpha}|H(x)|^{2}-2\langle\frac{D^{\Sigma}f(x)}{f},D^{\Sigma}u(x)\rangle-\langle H(x),y\rangle$
$\leq\log f(x)^{2}+\alpha|D^{\Sigma}u(x)|^{2}-\frac{1}{4\alpha}|H(x)|^{2}-\langle H(x),y\rangle$
$=\log f(x)^{2}+\alpha(|D^{\Sigma}u(x)|^{2}+y^{2})-\Big|\sqrt{\alpha}y+\frac{1}{2\sqrt{\alpha}}H(x)\Big|^{2},$
By Lemma 3.2 and the elementary inequality $\lambda\leq e^{\lambda-1}$, it follows that
$e^{-\alpha\frac{|d(x,\Phi_{r}(x,y))|^{2}}{r^{2}}}|\det D\Phi_{r}(x,y)|\leq r^{n+m}e^{\frac{n}{r}-n}e^{-\big|\sqrt{\alpha}y+\frac{1}{2\sqrt{\alpha}}H\big|^{2}}f(x)^{2}.$
Proof of Theorem 1.3.
By combining Lemma 3.1 and Lemma 3.3 with the formula for change of variables in multiple integrals, we obtain
$\int_{\Sigma_{r}}e^{-\alpha\frac{(r_{o}+d(o,\xi))^{2}}{r^{2}}}\,d\mathrm{vol}(\xi)$
$\leq\int_{\Sigma}\Big(\int_{T_{x}^{\perp}\Sigma}e^{-\alpha\frac{d(x,\Phi_{r}(x,y))^{2}}{r^{2}}}|\det D\Phi_{r}(x,y)|1_{A_{r}}(x,y)\,dy\Big)\,d\mathrm{vol}(x)$
$\leq r^{n+m}e^{\frac{n}{r}-n}\int_{\Sigma}f(x)^{2}\Big(\int_{T_{x}^{\perp}\Sigma}e^{-\big|\sqrt{\alpha}y+\frac{1}{2\sqrt{\alpha}}H(x)\big|^{2}}\,dy\Big)\,d\mathrm{vol}(x)$
$=r^{n+m}e^{\frac{n}{r}-n}\Big(\frac{\pi}{\alpha}\Big)^{\frac{m}{2}}\int_{\Sigma}f(x)^{2}\,d\mathrm{vol}(x).$
Note that $B_{o}\Big(\frac{r^{2}}{\alpha}-r_{o}\Big)\subset\Sigma_{r}$, the co-area formula implies
$\int_{0}^{\frac{r^{2}}{\alpha}-r_{o}}e^{-\alpha\frac{(r_{o}+\rho)^{2}}{r^{2}}}(n+m)\theta\omega_{n+m}\rho^{n+m-1}\,d\rho\leq r^{n+m}e^{\frac{n}{r}-n}\Big(\frac{\pi}{\alpha}\Big)^{\frac{m}{2}}\int_{\Sigma}f(x)^{2}\,d\mathrm{vol}(x).$
Dividing the above formula by $r^{n+m}$ and let $r\to\infty$, we have
$\theta e^{n}\Big(\frac{\pi}{\alpha}\Big)^{\frac{n}{2}}\leq\int_{\Sigma}f(x)^{2}\,d\mathrm{vol}(x).$
Consequently,
$n+\frac{n}{2}\log\frac{\pi}{\alpha}+\log\theta\leq\log\Big(\int_{\Sigma}f^{2}\,d\mathrm{vol}\Big).$
Combining this inequality with 4, one can get
$\int_{\Sigma}f^{2}\Big(\log f^{2}+n+\frac{n}{2}\log\frac{\pi}{\alpha}+\log\theta\Big)\,d\mathrm{vol}-\frac{1}{\alpha}\int_{\Sigma}|D^{\Sigma}f|^{2}\,d\mathrm{vol}$
$-\frac{1}{4\alpha}\int_{\Sigma}f^{2}|H|^{2}\,d\mathrm{vol}\leq\Big(\int_{\Sigma}f^{2}\,d\mathrm{vol}\Big)\log\Big(\int_{\Sigma}f^{2}\,d\mathrm{vol}\Big).$
Let $\int_{\Sigma}f^{2}\,d\text{vol}=1$, we have
$\int_{\Sigma}f^{2}\log f^{2}\,d\mathrm{vol}\leq-n-\frac{n}{2}\log\frac{\pi}{\alpha}-\log\theta+\frac{1}{\alpha}\int_{\Sigma}|Df|^{2}+\frac{1}{4}f^{2}|H|^{2}\,d\mathrm{vol}.$
Minimizing the right-hand side with respect to $\alpha>0$ yields
$\alpha=\frac{2}{n}\int_{\Sigma}|Df|^{2}+\frac{1}{4}f^{2}|H|^{2}\,d\mathrm{vol},$
which implies
$\int_{\Sigma}f^{2}\log f^{2}\leq\frac{n}{2}\log\Big(\frac{2}{\pi en}\theta^{-\frac{n}{2}}\int_{\Sigma}|Df|^{2}+\frac{1}{4}f^{2}|H|^{2}\,d\mathrm{vol}\Big).$